You are conducting a survey at a college with 800 students, 50 faculty members,
and 150 administrative staff members. Each of these 1,000 individuals has
a single listing in the campus phone directory. Suppose you were to cut up the
directory and pull out one listing at random to contact. What is the probability
it would be (a) a student, (b) a faculty member, (c) an administrative staff
member, (d) a faculty or administrative staff member, and (e) anyone except an
administrative staff member? (f) Explain your answers to someone who has
never had a course in statistics.

Answer by stanbon(75887)

useYou are conducting a survey at a college with 800 students, 50 faculty members,
and 150 administrative staff members. Each of these 1,000 individuals has
a single listing in the campus phone directory.
Comment: The numbers do not add up to 1000
-----------------------------------------------------
Suppose you were to cut up the directory and pull out one listing at random to contact.
What is the probability it would be
(a) a student,------800/1000 = 0.8
(b) a faculty member,------50/1000 = 1/20
(c) an administrative staff member,-----150/1000 = 0.15
(d) a faculty or administrative staff member,----(50+150)/1000 = 1/5((e) anyone except an administrative staff member?-----(1000-150)/1000 = 850/1000= 0.85
================================================
Probability = (# of ways to succeed)/(# of ways to succeed + # of ways to fail)
======================================================================
Cheers,
Stan H.
Reply to stanbon@comcast.net or use algebra.com response.